Solve for θ to the nearest hundredth, where 0 ≤ x ≤ 2π. a) 3 tan2 θ – 2 tan θ = 1


1
Expert's answer
2021-01-19T04:38:41-0500

Solution


  • First write the "\tan2\theta" in terms of "\tan\theta" to get the equation in like terms.
  • Then,

"\qquad\qquad
\begin{aligned}
3\bigg(\frac{2\tan\theta}{1-\tan^2\theta}\bigg)-2\tan\theta&=1\\
\end{aligned}"

  • For easy recognition, substitute "\tan\theta=x" & simplify the equation to find the roots of it.
  • Then,

"\qquad\qquad
\begin{aligned}
\frac{6x}{1-x^2}-2x&=0\\
\frac{2x^3+x^2+4x-1}{1-x^2}&=0
\end{aligned}"

  • Now "x \neq \pm1" for the equation to exist.
  • Then its about finding the roots of the numerator equation.
  • They are

"x=\begin{cases}
0.2306\\
-0.365+1.427i\\
-0.365-1.426i
\end{cases}"


  • All those values are acceptable only for an algebraic variable like x but only 0.2306 is valid for "\tan \theta" as "-\infty<\tan\theta<+\infty"
  • Therefore,

"\qquad\qquad
\begin{aligned}
x&=\tan\theta=0.2306\\
\theta&= 0.2266 rad
\end{aligned}"